OpenStudy (anonymous):

on a snowy day the amount of money Billy earns M varies directly with the number of driveways that he shovels D when he shovels 7 driveways he earns $80.50 how much dose he earn if he shovels 13 driveways $149.50 $11.50 $161.00 $113.00

OpenStudy (cggurumanjunath):

M=kD

OpenStudy (cggurumanjunath):

D = 7

OpenStudy (cggurumanjunath):

M = 80 . 50

OpenStudy (cggurumanjunath):

D = 13 M = k D

OpenStudy (anonymous):

149.5 is answer

OpenStudy (anonymous):

answer?

OpenStudy (cggurumanjunath):

M = k D 80.50 = k * 7 k = 80.50 / 7 = 11.5

OpenStudy (cggurumanjunath):

M = k D M = 11.5 * 13 = ? @samantha1995